# Two-Stage Pretraining for Molecular Property Prediction in the Wild

### v1

@@ Appendix A Appendix
A.1 Details of Datasets Used in the MPPW Benchmark
@@
The Molecular Property Prediction in the Wild (MPPW) benchmark uses two types of datasets: pretraining datasets and downstream datasets. For our first-stage pretraining, as well as in the pretraining of other models shown in Table 1, we randomly select 1M unlabeled molecules from the GDB17 dataset (Ruddigkeit et al., 2012). For the second-stage pretraining, we sample around 130K molecules from the 1M subset and calculate the auxiliary labels—HOMO, LUMO, and Dipole Moment—using Psi4 (Smith et al., 2020). This smaller subset is also used to pretrain other models shown in 8.
@@
For downstream evaluation, we curated 22 small datasets from the ChemBL database (Zdrazil et al., 2024), representing a diverse set of molecular properties as detailed in Table 7. To ensure consistency across datasets, we filter out any molecules containing atoms not present in the GDB17 dataset. As a result, only molecules containing the atoms {H, C, N, O, S, F, Cl, Br, I} are included in the downstream datasets. For evaluation, each dataset is randomly sampled to create three train/test splits with a 50:50 ratio, and all models in Tables 1 and 8 are assessed using these same splits. The processed datasets can be accessed through this URL.
@@
A.2 More results on the MPPW benchmark
As an additional experiment, we evaluated the downstream performance of MoleVers alongside the baseline models. In this experiment, the baselines are first pretrained using their original pretraining strategy, followed by a second-stage pretraining via auxiliary property prediction. The results, presented in Table 8, show that MoleVers achieves state-of-the-art MAE in 20 out of the 22 datasets, and ranks second in the remaining two. In terms of R2 scores, MoleVers achieves the best performance in 19 out of the 22 datasets, and ranks second in the other three. Note that none of the other models consistently rank in the top two across all datasets. Since all models are pretrained with the auxiliary labels, the results in Table 8 further highlight the benefits of our branching encoder which enables denoising pretraining with larger noise scales.
Table 8: Quantitative results on the MPPW benchmark. The ++ version of existing methods are trained in two stages: first with their vanilla pretraining strategy, then with auxiliary predictions. We show the mean MAE ($\downarrow$) and R2 ($\uparrow$) across three train/test splits. Bolded and underlined values are the best and second best results, respectively. The numbers within the parentheses after the assay ids are the number of training molecules in the assay.
GraphMVP++ MoleBert++ UniMol++ MolAE++ MoleVers
Assay MAE R2 MAE R2 MAE R2 MAE R2 MAE R2
1 (50) 0.482 0.803 0.567 0.720 0.460 0.806 0.446 0.759 0.448 0.820
2 (50) 0.375 0.386 0.297 0.524 0.303 0.534 0.356 0.306 0.281 0.582
3 (47) 3.512 -0.054 3.373 0.017 3.397 0.063 3.231 0.109 3.103 0.192
4 (24) 0.332 0.673 0.499 0.200 0.383 0.590 0.358 0.641 0.319 0.727
5 (26) 0.447 -0.263 0.590 -1.322 0.430 -0.377 0.431 -0.176 0.417 -0.246
6 (48) 26.624 -0.540 28.197 -0.494 18.54 -0.028 17.85 -0.013 17.66 -0.012
7 (35) 0.697 0.031 0.887 -0.493 0.723 -0.054 0.748 -0.179 0.630 0.179
8 (30) 0.645 -0.147 0.803 -0.829 0.687 -0.279 0.613 0.070 0.613 0.106
9 (34) 0.336 -0.250 0.330 0.020 0.319 0.107 0.330 0.090 0.312 0.170
10 (39) 0.218 0.335 0.196 0.348 0.226 0.313 0.220 0.348 0.187 0.490
11 (38) 0.426 0.043 0.451 0.020 0.448 0.050 0.432 0.049 0.413 0.171
12 (25) 0.738 0.251 0.738 0.256 0.653 0.337 0.653 0.327 0.611 0.412
13 (22) 0.627 0.089 0.627 -0.120 0.640 -0.125 0.604 0.053 0.484 0.329
14 (43) 0.343 -0.321 0.259 0.405 0.373 -1.119 0.310 0.084 0.280 0.314
15 (48) 0.426 0.619 0.441 0.547 0.572 0.199 0.467 0.519 0.385 0.665
16 (24) 0.928 0.124 0.719 0.360 0.838 0.231 0.839 0.076 0.700 0.364
17 (42) 1.186 0.141 1.284 -0.230 1.244 -0.062 1.274 -0.087 1.142 0.143
18 (31) 0.188 0.692 0.188 0.721 0.165 0.761 0.149 0.815 0.141 0.855
19 (62) 0.278 -0.360 0.250 -0.215 0.227 -0.126 0.192 0.160 0.191 0.259
20 (51) 0.289 -0.043 0.257 0.202 0.241 0.158 0.277 0.052 0.234 0.343
21 (19) 0.383 0.525 0.486 0.222 0.431 0.322 0.376 0.518 0.351 0.572
22 (22) 0.671 0.016 0.641 0.103 0.643 0.130 0.556 0.314 0.526 0.263
A.3 Impact of Finetuning Dataset Size on Downstream Performance

### v2

@@ Appendix A Appendix
A.1 Leveraging Large Language Models for Auxiliary Labels Generation
In this section, we discuss the use of large language models for generating auxiliary molecular property labels. In our experiments, we employ the ChatGPT 4o model as the label generator. The model is initialized with the prompt provided in Table 6, and a CSV file containing pairs of molecules along with the target property of interest are uploaded to the model. The predictions are saved in a downloadable CSV file. We have encountered instances where the model either declines to provide outputs or returns placeholder values (e.g., all 0s or 1s). In such cases, restarting the process from the initial prompt typically resolves the issue.
Table 6: Initial prompt to the LLM.
You will be given a csv file that consists of pairs of molecules represented as SMILES strings. The user will input the molecular property of interest in the same prompt. Your task is to (1) explain briefly what does the property mean, (2) explain briefly in a paragraph what makes a molecule exhibit a higher value in said property, (3) based on your explanation and existing knowledge, for each pair, rank the molecules. Output 0 if psmiles1 $>$ psmiles2, and 1 otherwise, where psmiles is the property of the molecule. Do not use any external tools. You should output your rank predictions as a csv file, where each line consists of ’smiles1, smiles2, prediction’.
There are two ways to utilize LLM-generated auxiliary labels: during the second pretraining stage or during finetuning. In MoleVers L2, we employ the LLM to predict pairwise rankings for 1 M pairs of molecules sampled from the GDB17 dataset. The molecular properties of the first seven assays in the MPPW benchmark serve as the target properties. In MoleVers D2L3, LLM predictions are incorporated during finetuning. First, for one assay, the LLM predicts pairwise rankings for all molecules in the test set. Then, the regression labels from the training split and the auxiliary ranking labels from the test split are used to finetune the model.
A.1.1 LLM as property ranker or regressor
The prompt in Table 6 can also be adapted to generate regression predictions, i.e., the absolute property values of each molecule. However, our experiments indicate that ChatGPT 4o performs better at predicting pairwise rankings than absolute property values. Table 7 compares the predictive performance of ChatGPT 4o for regression and ranking tasks across the first 13 assays in the MPPW benchmark. R2 and Kendall’s $\tau$ are employed as the evaluation metrics for regression and ranking, respectively. An ideal regression prediction would result in an $R^{2}$ score of 1, whereas simply predicting the mean would yield a score of 0. Similarly, a perfect ranking would correspond to a $|\tau|$ value of 1, while a completely randomized ranking would score 0. As shown in the table, ChatGPT consistently achieves equal or better performance in predicting pairwise rankings compared to absolute property values. Therefore, we use pairwise rankings as the auxiliary labels in our main experiments rather than absolute property values.
Table 7: Comparison of ChatGPT’s predictive performance for regression (absolute property values) and pairwise ranking (relative property values). We can see that ChatGPT consistently outputs higher-quality ranking predictions compared with absolute value predictions.
Assay ID 1 2 3 4 5 6 7 8 9 10 11 12 13
Regression ($R^{2}$) 0.50 0.22 0.00 0.37 0.00 0.01 0.15 0.10 0.05 0.05 0.02 0.14 0.00
Ranking ($|\tau|$) 0.50 0.34 0.04 0.44 0.06 0.08 0.17 0.11 0.08 0.24 0.05 0.23 0.13
A.1.2 Effects of LLM prediction quality to the downstream prediction performance
The performance of LLM ranking predictions can be evaluated using at least three metrics: accuracy, Kendall’s $\tau$, and the absolute value of Kendall’s $\tau$ ($|\tau|$). From our experiments, we observe weak correlations between MAE improvements and the first two metrics: -0.00557 for LLM accuracy and -0.00274 for LLM’s $\tau$. However, a much stronger correlation of -0.30563 is observed between MAE improvements and $|\tau|$. This indicates that there is a correlation between the quality of LLM ranking predictions and the downstream prediction performance (higher $|\tau|$ values leads to reduction in MAE).
It is noteworthy that this relationship is not evident when using accuracy or raw $\tau$ as metrics. This can be explained by the nature of the pairwise ranking, a binary classification task, employed during training with auxiliary labels. In this setting, LLM-predictions of very low accuracy (e.g., 0%) or $\tau$ score (e.g., -1) provides signals that are equally strong to very high accuracy (e.g., 100%) or $\tau$ score (e.g., +1).
To illustrate this, consider a set of LLM-generated labels with an accuracy of 50%. Such a set is equivalent to random noise, offering no useful information. However, a set of labels with 0% accuracy, while seemingly poor by conventional metrics, contains perfect (although inverted) information, as every prediction is consistently wrong. Similarly, a $\tau$ score of -1 indicates a perfect reverse ranking, holding the same signal strength as a $\tau$ score of +1. For this reason, the absolute value of Kendall’s $\tau$ ($|\tau|$) is a more reliable predictor of downstream MAE improvements compared to raw accuracy or $\tau$ scores.
A.1.3 In-context learning for LLM ranking
One approach to improve a large language model on a given task is by providing it with relevant examples, a process also known as in-context learning. Here, we explore the possibility of adapting ChatGPT 4o to our assays through in-context learning. For a given assay, we provide the model with molecules from the training split, along with their corresponding property labels. The model is then tasked with generating the auxiliary ranking labels for the test molecules.
As shown in the box plots in Figure 5, incorporating additional data generally improves both accuracy and $\tau$ values. However, the absolute $\tau$ values remain relatively unchanged. This is because the observed improvements primarily stem from re-aligning the LLM predictions to the pre-determined labeling convention. For example, ChatGPT might yield a 0-shot $\tau$ value of -0.26 and an in-context $\tau$ value of 0.09. In this scenario, the 0-shot model tends to predict flipped rankings correctly, while the in-context model more accurately predicts the actual rankings. Although the in-context approach shifts the $\tau$ value in a positive direction, the magnitude is deacreasing. As a result, the in-context model shows improvement in overall $\tau$, but not in absolute $|\tau|$.
Figure 5: Predictive performance comparison between ChatGPT 0-shot and ChatGPT with in-context learning across 22 assays in the MPPW benchmark. While in-context learning improves the predictive accuracy and Kendall’s tau in general, the absolute value of Kendall’s tau remains relatively unchanged.
A.2 Network and Training Details
The primary and auxiliary encoders of MoleVers are built on the UniMol encoder architecture (Zhou et al., 2023). Each encoder comprises 15 layers, with an embedding dimension of 512 and a feedforward dimension of 2048. The MAP and denoising heads are implemented with multilayer perceptrons, while the aggregator module is implemented with pooling by multihead attention (PMA), a cross attention-based module introduced by Lee et al. (2019). The PMA module uses a query of size $1\times 512$, and takes the molecule features $F$ as key and value. During the first pretraining stage, the model is trained for 1 million iterations using a batch size of 32, with a masking ratio of 0.15 for the MAP task. In the second pretraining stage, the model is trained for 50 epoch, maintaining the same batch size of 32. We employ the Adam optimizer with a learning rate of 10-4 and utilize a polynomial decay learning rate scheduler. We run all experiments on an NVIDIA Quadro RTX 8000 GPU.
A.3 Details of datasets used in of the MPPW Benchmark
@@
The Molecular Property Prediction in the Wild (MPPW) benchmark uses two types of datasets: pretraining datasets and downstream datasets. For our first-stage pretraining, as well as in the pretraining of other models shown in Table 9, we randomly select 1M unlabeled molecules from the GDB17 dataset (Ruddigkeit et al., 2012). For the second-stage pretraining, we sample around 130K molecules from the 1M subset and calculate the auxiliary labels—HOMO, LUMO, and Dipole Moment—using Psi4 (Smith et al., 2020).
@@
For downstream evaluation, we curated 22 small datasets from the ChemBL database (Zdrazil et al., 2024), representing a diverse set of molecular properties as detailed in Table 8. To ensure consistency across datasets, we filter out any molecules containing atoms not present in the GDB17 dataset. As a result, only molecules containing the atoms {H, C, N, O, S, F, Cl, Br, I} are included in the downstream datasets. For evaluation, each dataset is randomly sampled to create three train/test splits with a 50:50 ratio, and all models in Tables 9 are assessed using these same splits. The processed datasets can be accessed through this URL.
@@
A.4 Details of experimental results in the MPPW benchmark
Table 9 presents the numerical values of the metrics discussed in Section 5.2. In addition to the MAE and R2 scores, we report the Kendall rank correlation coefficient ($\tau$) for MoleVers and the baselines. Kendall’s $\tau$ ranges from -1 to 1, where a perfect rank prediction yields +1, a random ranking results in 0, and a score of -1 indicates a completely inverted ranking relative to the ground truth. The rank correlation coefficient is particularly relevant since performance optimization is a common objective in real-world applications Identifying molecules with better properties (e.g., higher solubility, higher activity), is therefore as important as predicting the absolute property values. To illustrate the results, we provide box plots of $\tau$ scores in Figure 6. Similar to the MAE and R2 results, MoleVers outperforms all baselines by a substantial margin.
Figure 6: Box plot of Kendall’s tau rank correlation coefficient (absolute values, $\uparrow$). The whiskers extend from the box to the farthest data point lying within 1.5x the inter-quartile range from the box.
Table 9: Quantitative results on the MPPW benchmark. We report the mean MAE ($\downarrow$), R2 ($\uparrow$), and Kendall’s Tau-b across 3 train/test splits for each assay.
Assay ID GraphMVP MoleBert UniMol
MAE R2 $\tau$ MAE R2 $\tau$ MAE R2 $\tau$
1 0.426 0.846 0.763 0.526 0.724 0.668 0.459 0.802 0.764
2 0.408 0.260 0.398 0.303 0.528 0.617 0.455 0.067 0.267
3 3.459 0.024 0.229 3.351 -0.011 0.296 3.282 0.035 0.252
4 0.329 0.716 0.725 0.429 0.500 0.619 0.522 0.186 0.573
5 0.463 -0.201 -0.093 2.277 -19.732 0.251 0.415 -0.129 0.391
6 0.269 -0.653 0.053 0.197 -0.066 0.070 0.180 -0.018 0.016
7 0.660 0.169 0.292 0.522 0.421 0.498 0.672 -0.060 0.251
8 0.810 -0.829 -0.205 3.449 -19.905 0.279 0.667 0.017 0.204
9 0.341 0.102 0.299 0.261 0.328 0.407 0.347 -0.011 0.133
10 0.238 0.285 0.398 0.418 -0.693 0.535 0.220 0.300 0.437
11 0.409 0.178 0.347 0.456 -0.107 0.258 0.428 0.019 0.362
12 0.727 0.133 0.396 0.797 0.149 0.398 0.653 0.316 0.582
13 0.630 0.051 0.147 1.148 -1.969 0.023 0.592 0.068 0.308
14 0.343 0.036 0.334 0.443 -0.351 0.430 0.330 0.132 0.381
15 0.378 0.703 0.627 0.563 0.356 0.500 0.474 0.512 0.586
16 0.753 0.357 0.455 0.784 0.351 0.471 0.946 -0.039 0.355
17 1.099 0.238 0.373 1.151 0.023 0.351 1.347 -0.390 0.220
18 0.172 0.754 0.765 0.163 0.815 0.790 0.181 0.719 0.804
19 0.305 -0.485 -0.102 0.307 -0.687 0.321 0.223 0.012 0.179
20 0.329 -0.402 0.067 0.271 0.089 0.362 0.303 0.000 0.217
21 0.343 0.622 0.680 0.375 0.532 0.583 0.589 -0.123 0.532
22 0.805 0.148 0.247 5.086 -23.570 0.600 0.616 0.251 0.438
Assay ID MolAE MoleVers
MAE R2 $\tau$ MAE R2 $\tau$
1 0.607 0.698 0.787 0.393 0.832 0.770
2 0.356 0.305 0.502 0.298 0.516 0.554
3 3.470 -0.096 0.239 2.999 0.211 0.370
4 0.475 0.399 0.706 0.332 0.701 0.740
5 0.402 0.012 0.320 0.401 -0.134 0.336
6 0.180 -0.016 0.057 0.179 -0.016 0.059
7 0.742 -0.169 0.158 0.650 0.018 0.252
8 0.630 0.079 0.249 0.593 0.144 0.310
9 0.351 -0.042 0.194 0.339 0.034 0.220
10 0.286 -0.089 0.313 0.190 0.457 0.481
11 0.444 -0.110 0.312 0.372 0.302 0.432
12 0.649 0.425 0.560 0.536 0.538 0.598
13 0.611 0.156 0.358 0.522 0.181 0.335
14 0.328 0.093 0.368 0.295 0.287 0.454
15 0.477 0.522 0.572 0.447 0.591 0.546
16 0.874 0.058 0.451 0.748 0.278 0.557
17 1.451 -0.533 0.162 1.078 0.192 0.254
18 0.227 0.572 0.592 0.152 0.822 0.795
19 0.230 -0.130 0.243 0.220 -0.188 0.365
20 0.296 -0.046 0.203 0.235 0.347 0.220
21 0.513 0.132 0.407 0.326 0.648 0.684
22 0.654 0.068 0.214 0.519 0.369 0.463
A.5 Results on Large Downstream Datasets
We further evaluate the performance of MoleVers on the MoleculeNet benchmark (Wu et al., 2018), focusing on large-scale regression datasets such as QM7, QM8, and QM9. As shown in Table 10, MoleVers outperforms all baseline models across all datasets, achieving the lowest MAE scores. Therefore, the proposed two-stage pretraining framework is not only effective in low-data regimes, but also excels when abundant labeled data is available. We note, however, scenarios in which thousands of labeled data is available is extremely rare in the real-world.
. Dataset QM7 QM8 QM9 #Molecules 6830 21789 133885 D-MPNN 103.5 0.0190 0.0081 Attentive FP 72.0 0.0179 0.0081 Pretrain-GNN 113.2 0.0200 0.0092 GROVER 94.5 0.0218 0.0099 MolCLR 66.8 0.0178 - Uni-Mol 58.9 0.0160 0.0054 Mol-AE 53.8 0.0161 0.0053 MoleVers (ours) 51.3 0.0155 0.0050
A.6 Impact of Finetuning Dataset Size on Downstream Performance
@@
Figure 8: Comparison of noise scale distributions when sampled using static vs. dynamic noise scales. The dynamic noise scale sampling biases the distribution towards noise scale of lower values, but still maintain a good amount of noise scale of higher values.